# Mixture of Directed Graphical Models for Discrete Spatial Random Fields

### v3

@@ Abstract
Current approaches for modeling discrete-valued outcomes associated with spatially-dependent areal units incur computational and theoretical challenges, especially in the Bayesian setting when full posterior inference is desired. As an alternative, we propose a novel statistical modeling framework for this data setting, namely a mixture of directed graphical models (MDGMs). The components of the mixture, directed graphical models, can be represented by directed acyclic graphs (DAGs) and are computationally quick to evaluate. The DAGs representing the mixture components are selected to correspond to an undirected graphical representation of an assumed spatial contiguity/dependence structure of the areal units, which underlies the specification of traditional modeling approaches for discrete spatial processes such as Markov random fields (MRFs). We introduce the concept of compatibility to show how an undirected graph can be used as a template for the structural dependencies between areal units to create sets of DAGs which, as a collection, preserve the structural dependencies represented in the template undirected graph. We then introduce three classes of compatible DAGs and corresponding algorithms for fitting MDGMs based on these classes. In addition, we compare MDGMs to MRFs and a popular Bayesian MRF model approximation used in high-dimensional settings in a series of simulations and an analysis of ecometrics data collected as part of the Adolescent Health and Development in Context Study.
Keywords: Adolescent Health and Development in Context study, areal data, Bayesian statistics, graph theory, Markov random field, spanning trees, spatial statistics
@@ 1 Introduction
Statistical models for spatially-dependent, discrete-valued observations associated with areal units (i.e., discrete space) are used routinely in areas of application ranging from disease mapping to small area estimation to image analysis. [^2] As “discrete” is used to describe multiple features of the analytic setting, we standardize our use of the term. We use discrete-valued observations/latent variables to refer to the observations/latent variables associated with areal units that take on discrete values. We use discrete spatial process to refer generally to the areal unit/lattice setting. Despite a rich literature on both classical and Bayesian models for this setting, theoretical and computational considerations remain, particularly in the Bayesian setting when complete posterior inference is desired. To address the computational and theoretical concerns of existing models (described below), we propose a novel framework for modeling spatial dependence of observations associated with a collection of $n$ areal units which together form a nonoverlapping partition of the spatial/geographic domain. The areal units may be a regular grid/lattice or be an irregular partition (e.g., administrative districts). Without loss of generality and to motivate our new framework, we assume at each areal unit there is a collection of zero-one binary observations, $[y_{i1},\dots,y_{im_{i}}]$, where $m_{i}$ is the number of observations at areal unit $i$. Additionally, we assume that there is a single binary latent variable, $z_{i}$, associated with each areal unit for $i=1,\dots,n$. Let $\mathbf{y}=[[y_{11},\dots,y_{1m_{1}}],\dots,[y_{n1},\dots,y_{nm_{n}}]]$ denote the complete collection of observations and $\mathbf{z}=[z_{1},\dots,z_{n}]$ the vector of binary latent variables.
@@
Standard spatial statistical models for areal data assume that there is a known natural undirected graph (NUG, our terminology) with vertices/edges denoting the areal units and the spatial proximity/contiguity of pairs of areal units, respectively. Markov random fields (MRFs) are a default choice for prior distribution on $\mathbf{z}$, as an MRF, defined for a given NUG, is in fact a probability distribution whose conditional distributions elicit the same dependence structure (i.e. dependencies between areal units) as the specified NUG (Besag, 1974, 1975). While the one-to-one correspondence between a NUG and an MRF leads to a simple and intuitive specification of the joint probability distribution of $\mathbf{z}$, inference on the parameters governing an MRF is computationally burdensome due to an intractable normalizing constant when $\mathbf{z}$ is discrete. While diverse approximation methods for the MRF have been proposed (Reeves and Pettitt, 2004; Friel et al., 2009; McGrory et al., 2009, 2012), the approach that uses the pseudo-likelihood (PL) of Besag (1975) in place of the analytical form of the intractable MRF prior density persists in Bayesian analyses (Heikkinen and Hogmander, 1994; Hoeting et al., 2000; Wang et al., 2000; Hughes et al., 2011; Pereyra and McLaughlin, 2017; Marsman and Haslbeck, 2023). While relatively easy to implement and computationally efficient, this approach lacks rigor as there are no guarantees that the approximation leads to valid posterior inference (see Section 3.3).
@@
As an alternative to specifying a joint probability model for the observed and latent variables associated with the areal units based on the assumed NUG directly, we propose using the NUG as a template for the structural dependencies that exist between the areal units. From the NUG we derive a collection of directed acyclic graphs (DAGs) which is used to specify probability distributions, directed graphical models (DGMs), that together capture the spatial dependence between random variables associated with the areal units. [^3] Directed graphical models are also called Bayesian networks in older literature, a name we dislike as there is nothing inherently Bayesian about a directed graphical model. The main advantage of a DGM is the straightforward factorization of the probability distribution into conditional distributions, which leads to quick computational evaluation of a DGM (Section 3). Let $\mathcal{D}(\mathbf{N})=\{\mathbf{D}_{1},\dots,\mathbf{D}_{k}\}$ be a set of DAGs which honor the structural dependencies of the NUG, $\mathbf{N}$, where $\mathbf{N}$ and $\mathbf{D}$ define the structure of a NUG and DAG, respectively, through a set of vertices and edges; we give formal definitions of $\mathbf{N}$ and $\mathbf{D}$ in Section 2. In order to formally describe how a DAG can be derived from a template NUG which honors the topology of the NUG, we define the notion of compatibility in Section 2.1. Given the set of DAGs, $\mathcal{D}$, and a DGM for $\mathbf{z}$, written as $p(\mathbf{z}|\mathbf{D},\boldsymbol{\xi})$, that depends on the structure of the DAG and parameters $\boldsymbol{\xi}$, we specify the prior on $\mathbf{z}$ as
@@
a mixture of directed graphical models (MDGM), where the mixture weights are given implicity by the prior over the collection of DAGs, $p(\mathbf{D})$ for $\mathbf{D}\in\mathcal{D}$. Given the latent $\mathbf{z}$, we model each $y_{ij}$ as conditionally independent Bernoulli variables,
@@
with noise parameters $\boldsymbol{\eta}=[\eta_{0},\eta_{1}]$. Let $\eta_{0}$ be the probability that $y_{i}=1$ when $z_{i}=0$ and $\eta_{1}$ be the probability that $y_{i}=1$ when $z_{i}=1$. In this setting the $y_{i}$’s can be viewed as a noisy version of the true underlying black and white image $\mathbf{z}$. For identifiability of the noise parameters, $\boldsymbol{\eta}$, and DGM parameters, $\boldsymbol{\xi}$, in the model we require $m_{i}>1$ for a subset of the areal units. While continuous-valued distributions for $\mathbf{y}$ given the latent $\mathbf{z}$ are more common in the literature, we focus on this discrete outcome setting with spatial dependence introduced through a discrete latent variable as fully Bayesian approaches in this setting have been less explored (Arnesen and Tjelmeland, 2015). We note, however, the ideas presented in the MDGM framework are generally applicable to the variety of combinations of discrete or continuous latent and observed variables. Additionally, while we present the MDGM as a prior in a hierarchical framework, the MDGM can also serve as a direct data model for discrete-valued outcomes associated with areal units.
The main contributions of the paper are as follows. In Section 2, we define the notion of compatibility, which provides a principled way to construct a DAG (and the corresponding DGM) from a template NUG. Through compatibility, sets of DAGs can be specified which are distinguished by some feature of the DAGs in the set, e.g. minimally connected DAGs, namely spanning trees, or fully connected DAGs, namely acyclic orientations of the NUG. We explore spanning trees, acyclic orientations as well a third class of DAGs, providing MCMC algorithms for each in Section 4. Compatibility expands the statistical toolkit by linking relevant graph theory to statistical modeling and by providing a formal foundation for the development of new classes of DAGs to be used in the MDGM, as specified in Section 3. Our notion of compatibility unifies existing approaches which have sought to utilize DAGs and other graph structures to facilitate inference, though often in a different inferential setting such as structure learning (Wu and Doerschuk, 1995; Meila and Jordan, 1997, 2001; Pletscher et al., 2009; Thiesson et al., 1999). We compare our models, via simulation and data analysis (Sections 5 and 6), using an MDGM prior to corresponding models with an MRF prior. In addition, we include a comparison to the widely-used, but theoretically ungrounded, “approximate MRF prior” (aMRF), discussed in Section 3.3.
@@ 2 The NUG as a Template for a DAG
In this section, we introduce the concept of compatibility, which is the criteria used to determine whether a DAG or set of DAGs retain structural integrity to the NUG. NUGs and DAGs are each graphical representations of the probability models, MRFs and DGMs respectively. We address specification of an MRF given a NUG and a DGM given a DAG in Section 3; for now, we focus on how the topology of the NUG serves as a template for the structural dependencies between vertices. To this end we introduce some notation.
Without loss of generality, we focus discussion on the NUG as the template for sets of DAGs, though the concepts apply to any undirected graph. Let $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, be a NUG with set of vertices, $\mathbf{V}=\{i:i=1,\dots,n\}$, where $n$ is the number of areal units, and set of unordered pairs, $\bar{\mathbf{E}}$, where the unordered pair, $\{i,j\}$, denotes that there is an undirected edge between vertices $i$ and $j$. In a NUG, we say that vertices $i$ and $j$ are neighbors if $\{i,j\}$ is an element of $\bar{\mathbf{E}}$. We introduce the notation
@@
A key feature of the NUG, is that the undirected edge, $\{i,j\}$, represents a symmetric relationship between vertices $i$ and $j$. In contrast, a DAG, defined on the same set of vertices and denoted $\mathbf{D}=\{\mathbf{V},\vec{\mathbf{E}}\}$, has an edge set, $\vec{\mathbf{E}}$, comprised of ordered pairs of vertices $(i,j)$, which we use to indicate, somewhat unconventionally, that there is a directed edge from $j$ to $i$. For the directed edge $(i,j)$ we say that $j$ is a parent of $i$ and define
@@
2.1 Compatibility of a graph to a NUG
How the NUG serves as a template of structural dependencies becomes clear with an alternative representation of the graph structure. Let $A(\mathbf{G})$ be an matrix representation of the graph $\mathbf{G}$. This association matrix, $A(\mathbf{G})$, is constructed by setting the $(i,j)$ and $(j,i)$ positions of the matrix to one if the undirected edge, $\{i,j\}$, is in the edge set and zero otherwise. For directed edges, $(i,j)$, in the edge set, only the $(i,j)$ position is set to one. By definition, the matrix $A(\mathbf{N})$ is symmetric. Now we can define the compatibility of a graph with the NUG.
@@ 2.1 Compatibility of a graph to a NUG
A graph, $\mathbf{G}=\{\mathbf{V},\mathbf{E}\}$, with vertices, $\mathbf{V}$, is compatible with a NUG, $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, defined on the same set of vertices, denoted $\mathbf{G}\Bumpeq\mathbf{N}$, when $(i,j)$ is equal to one in $A(\mathbf{G})$ implies that $(i,j)$ is equal to one in $A(\mathbf{N})$.

### v4

@@ Abstract
Current approaches for modeling discrete-valued outcomes associated with spatially-dependent areal units incur computational and theoretical challenges, especially in the Bayesian setting when full posterior inference is desired. As an alternative, we propose a novel statistical modeling framework for this data setting, namely a mixture of directed graphical models (MDGMs). The components of the mixture, directed graphical models, can be represented by directed acyclic graphs (DAGs) and are computationally quick to evaluate. The DAGs representing the mixture components are selected to correspond to an undirected graphical representation of an assumed spatial contiguity/dependence structure of the areal units, which underlies the specification of traditional modeling approaches for discrete spatial processes such as Markov random fields (MRFs). Notably, the MDGM is not proposed as an approximation to an MRF, but rather shares the same default, underlying graphical representation of spatial dependence as an MRF. However, in the case that the data generating mechanism of the latent spatial field is known to be an MRF, we find that posterior inferences under an MDGM prior better approximate the posterior of the model with a correctly specified MRF prior. We introduce the concept of compatibility to show how an undirected graph can be used as a template for the dependencies between areal units to create sets of DAGs which, as a collection, preserve the dependencies represented in the template undirected graph. Lastly, we compare highlighted classes of MDGMs to MRFs and a popular Bayesian MRF model approximation used in high-dimensional settings in a series of simulations and an analysis of ecometrics data collected as part of the Adolescent Health and Development in Context Study.
Keywords: Adolescent Health and Development in Context study, areal data, Bayesian statistics, graph theory, Markov random fields, spanning trees, spatial statistics
@@ 1 Introduction
Statistical models for spatially-dependent, discrete-valued observations associated with areal units (i.e., discrete space) are used routinely in areas of application ranging from disease mapping to small area estimation to image analysis. As “discrete” is used to describe multiple features of the analytic setting, we standardize our use of the term. We use discrete-valued observations/latent variables to refer to the observations/latent variables associated with areal units that take on discrete values. We use discrete spatial process to refer generally to the areal unit/lattice setting. Despite a rich literature on both classical and Bayesian models for this setting, theoretical and computational considerations remain, particularly in the Bayesian setting when complete posterior inference is desired. To address the computational and theoretical concerns of existing models (described below), we propose a novel framework for modeling spatial dependence of observations associated with a collection of $n$ areal units which together form a nonoverlapping partition of the spatial/geographic domain. The areal units may be a regular grid/lattice or be an irregular partition (e.g., administrative districts). Without loss of generality and to motivate our new framework, we assume at each areal unit there is a collection of zero-one binary observations, $[y_{i1},\dots,y_{im_{i}}]$, where $m_{i}$ is the number of observations at areal unit $i$. Additionally, we assume that there is a single binary latent variable, $z_{i}$, associated with each areal unit for $i=1,\dots,n$. Let $\mathbf{y}=[[y_{11},\dots,y_{1m_{1}}],\dots,[y_{n1},\dots,y_{nm_{n}}]]$ denote the complete collection of observations and $\mathbf{z}=[z_{1},\dots,z_{n}]$ the vector of binary latent variables.
@@
Standard spatial statistical models for areal data assume that there is a known natural undirected graph (NUG, our terminology) with vertices/edges denoting the areal units and the spatial proximity/contiguity of pairs of areal units, respectively. Markov random fields (MRFs) are a default choice for prior distribution on $\mathbf{z}$, as an MRF, defined for a given NUG, is in fact a probability distribution whose conditional distributions elicit the same dependence structure (i.e. dependencies between areal units) as the specified NUG (Besag, 1974, 1975). While the one-to-one correspondence between a NUG and an MRF leads to a simple and intuitive specification of the joint probability distribution of $\mathbf{z}$, inference on the parameters governing an MRF is computationally burdensome due to an intractable normalizing constant when $\mathbf{z}$ is discrete. While diverse approximation methods for the MRF have been proposed (Reeves and Pettitt, 2004; Friel et al., 2009; McGrory et al., 2009, 2012), the approach that uses the pseudo-likelihood of Besag (1975) in place of the analytical form of the intractable MRF prior density persists in Bayesian analyses (Heikkinen and Hogmander, 1994; Hoeting et al., 2000; Wang et al., 2000; Hughes et al., 2011; Pereyra and McLaughlin, 2017; Marsman and Haslbeck, 2023). While relatively easy to implement and computationally efficient, this approach lacks rigor as there are no guarantees that the approximation leads to valid posterior inference (see Section 3.3).
@@
As an alternative to specifying a joint probability model for the observed and latent variables associated with the areal units based on the assumed NUG directly, we propose using the NUG as a template for the dependencies that exist between the areal units. From the NUG, we derive a collection of directed acyclic graphs (DAGs) which is used to specify probability distributions, directed graphical models (DGMs), that together capture the spatial dependence between random variables associated with the areal units. Directed graphical models are also called Bayesian networks in older literature, a name we dislike as there is nothing inherently Bayesian about a directed graphical model. The main advantage of a DGM is the straightforward factorization of the probability distribution into conditional distributions, which leads to quick computational evaluation of a DGM (Section 3). Let $\mathcal{D}(\mathbf{N})=\{\mathbf{D}_{1},\dots,\mathbf{D}_{k}\}$ be a set of DAGs which honor the dependencies of the NUG, $\mathbf{N}$, where $\mathbf{N}$ and $\mathbf{D}$ define the structure of a NUG and DAG, respectively, through a set of vertices and edges; we give formal definitions of $\mathbf{N}$ and $\mathbf{D}$ in Section 2. In order to formally describe how a DAG can be derived from a template NUG which honors the topology of the NUG, we define the notion of weak compatibility in Section 2.1. In Section 2.2, we provide a more robust definition of compatibility, which also ensures fidelity of the DAG to the conditional independence relationships of the NUG. Given the set of DAGs, $\mathcal{D}$, and a DGM for $\mathbf{z}$, written as $p(\mathbf{z}|\mathbf{D},\boldsymbol{\xi})$, that depends on the structure of the DAG and parameters $\boldsymbol{\xi}$, we specify the prior on $\mathbf{z}$ as
@@
a mixture of directed graphical models (MDGM), where the mixture weights are given implicity by the prior over the collection of DAGs, $p(\mathbf{D})$ for $\mathbf{D}\in\mathcal{D}$. Given the latent $\mathbf{z}$, we model each $y_{ij}$ as conditionally independent Bernoulli random variables,
@@
with noise parameters $\boldsymbol{\eta}=[\eta_{0},\eta_{1}]$, where $\eta_{0}$ is the probability that $y_{i}=1$ when $z_{i}=0$ and $\eta_{1}$ is the probability that $y_{i}=1$ when $z_{i}=1$. In this setting the $y_{i}$’s can be viewed as a noisy version of the true underlying binary image $\mathbf{z}$. A realization $\mathbf{z}$ of the spatial field with respect to the NUG is called a configuration of the graph. We call realizations from both MRFs and MDGMs configurations of the NUG. While continuous-valued distributions for $\mathbf{y}$ given the latent $\mathbf{z}$ are more common in the literature, we focus on this discrete outcome setting with spatial dependence introduced through a discrete latent variable as fully Bayesian approaches in this setting have been less explored (Arnesen and Tjelmeland, 2015). We note, however, the ideas presented in the MDGM framework are generally applicable to the variety of combinations of discrete or continuous latent and observed variables. Additionally, while we present the MDGM as a prior in a hierarchical framework, the MDGM can also serve as a direct data model for discrete-valued outcomes associated with areal units.
The main contributions of the paper are as follows. In Section 2, we define the notion of compatibility, with its weak and strong versions, which provides a principled way to construct a DAG (and the corresponding DGM) from a template NUG. Through compatibility, sets of DAGs can be specified which are distinguished by some feature of the DAGs in the set, e.g. minimally connected DAGs, namely spanning trees, or fully connected DAGs, namely acyclic orientations of the NUG. We highlight the class of spanning trees, which are strongly compatible with any NUG and discuss model fitting via MCMC in Section 4. Additional MDGM classes are presented in the Appendix. Compatibility expands the statistical toolkit by linking relevant graph theory to statistical modeling and by providing a formal foundation for the development of new classes of DAGs to be used in the MDGM, as specified in Section 3. Our notion of compatibility unifies existing approaches which have sought to utilize DAGs and other graph structures to facilitate inference, though often in a different inferential setting such as structure learning (Wu and Doerschuk, 1995; Meila and Jordan, 1997, 2001; Pletscher et al., 2009; Thiesson et al., 1999). We compare our models, via simulation and data analysis (Sections 5 and 6), using an MDGM prior to corresponding models with an MRF prior. In addition, we include a comparison to the widely-used, but theoretically ungrounded, “approximate MRF prior” (aMRF), discussed in Section 3.3.
@@ 2 The NUG as a Template for a DAG
In this section, we introduce the concept of compatibility, which is the criteria used to determine whether a DAG or set of DAGs honor the topology of the NUG. DAGs which honor the structure of the NUG are said to be weakly compatible, while DAGs which honor the dependence relationships implied by the topology of the NUG are strongly compatible. We formally define weak and strong compatibility below.
NUGs and DAGs are each graphical representations of probability models, MRFs and DGMs, respectively. We address specification of an MRF given a NUG and a DGM given a DAG in Section 3; for now, we begin with weak compatibility and focus on how the topology of the NUG serves as a template for the structural dependencies between vertices. To this end we introduce some notation.
Without loss of generality, we center discussion on the NUG as the template for sets of DAGs, though the concepts apply to any undirected graph. Let $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, be a NUG with set of vertices, $\mathbf{V}=\{i:i=1,\dots,n\}$, where $n$ is the number of areal units, and set of unordered pairs, $\bar{\mathbf{E}}$, where the unordered pair, $\{i,j\}$, denotes that there is an undirected edge between vertices $i$ and $j$. In a NUG, we say that vertices $i$ and $j$ are neighbors if $\{i,j\}$ is an element of $\bar{\mathbf{E}}$. We introduce the notation
@@
A key feature of the NUG is that the undirected edge, $\{i,j\}$, represents a symmetric relationship between vertices $i$ and $j$. In contrast, a DAG, defined on the same set of vertices and denoted $\mathbf{D}=\{\mathbf{V},\vec{\mathbf{E}}\}$, has an edge set, $\vec{\mathbf{E}}$, comprised of ordered pairs of vertices $(i,j)$, which we use to indicate, somewhat unconventionally, that there is a directed edge from $j$ to $i$. For the directed edge $(i,j)$ we say that $j$ is a parent of $i$ and define
@@
2.1 Weak compatibility of a graph to a NUG
How the NUG serves as a template of structural dependencies becomes clear with an alternative representation of the graph structure. Let $A(\mathbf{G})$ be a matrix representation of the graph $\mathbf{G}$. This adjacency matrix, $A(\mathbf{G})$, is constructed by setting the $(i,j)$ and $(j,i)$ positions of the matrix to one if the undirected edge, $\{i,j\}$, is in the edge set and zero otherwise. For directed edges, $(i,j)$, in the edge set, only the $(i,j)$ position is set to one. By definition, the matrix $A(\mathbf{N})$ is symmetric. A property of the adjacency matrix for a DAG, $\mathbf{D}$, is that there exists a permutation of the rows and columns of $A(\mathbf{D})$ for which $A(\mathbf{D})$ is lower triangular. Additionally, in a DAG, the $i$th row of $A(\mathbf{D})$ is a $m$-dimensional vector with binary indicators for the parents of $i$, that is $\pi(i)$. Now we can define weak compatibility of a graph with the NUG.
@@ 2.1 Weak compatibility of a graph to a NUG
A graph, $\mathbf{G}=\{\mathbf{V},\mathbf{E}\}$, with vertices, $\mathbf{V}$, is weakly compatible with a NUG, $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, defined on the same set of vertices, denoted $\mathbf{G}\bumpeq\mathbf{N}$, when $(i,j)$ is equal to one in $A(\mathbf{G})$ implies that $(i,j)$ is equal to one in $A(\mathbf{N})$.